OpenStudy (anonymous):

Find an equation of the circle that satisfies the given conditions. Center (−4, 7); passes through (−9, −8)

OpenStudy (whpalmer4):

Formula for a circle with radius \(r\) and center \((h,k)\) is \[(x-h)^2 + (y-k)^2 = r^2\] Formula for the distance between two points \((x_1,y_1)\) and \((x_2,y_2)\) is \[d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\] Use the second formula to find the radius, which is simply the distance between the center and a point through which the circle passes. With the radius in hand, fill in the blanks in the formula for the circle and you're done.

OpenStudy (whpalmer4):

Because the circle formula asks for \(r^2\), you don't have to even both with the square root in the distance formula: \(r^2 = (x_2-x_1)^2 + (y_2-y_1)^2\)

OpenStudy (anonymous):

Show that the equation represents a circle by rewriting it in standard form. 4x2 + 4y2 + 8x − y = 0

OpenStudy (whpalmer4):

Complete the square on x and y to achieve that.

OpenStudy (anonymous):

but how when it equals zero? i get that i would have to divide everything by 4 but then i will never have a radius